OpenStudy (anonymous):

Mike is using a telescope to view three stars. He notes the distances and angles between the lines joining the stars, as shown below. Which expression shows the distance between Star 2 and Star 3?

OpenStudy (anonymous):

A) y sec x° B) y over sec x° C) y cot x° D) y cot over x°

OpenStudy (anonymous):

cosx = y/d, where d is the distance you want... but this can also be written as secx = d/y and if you solve for d in that second equation, you'll have your answer... :)

OpenStudy (anonymous):

could you like step by step explain it

OpenStudy (anonymous):

sure... do you agree that in the right triangle in the picture, cosx = y/d ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok... secx = 1/cosx, do you agree with this? it is a basic identity....

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

another way to say that ^^ is cosx = 1/secx

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok so...

OpenStudy (anonymous):

wait on the cosx thing why is the x there?

OpenStudy (anonymous):

x is the angle...

OpenStudy (anonymous):

ok i think i got it but it will take a min

OpenStudy (anonymous):